OpenStudy (anonymous):

State the various transformations applied to the base function ƒ(x) = |x| to obtain a graph of the function g(x) = 3[|x − 1| + 2]

OpenStudy (p0sitr0n):

first there is a vertical stretch of 3 then you translate the function 2 units up lastly you move the function horizontally 1 to the right

OpenStudy (p0sitr0n):



OpenStudy (anonymous):

what would the equation for this be http://tinyurl.com/kmssyap

OpenStudy (p0sitr0n):

i would guess g(x)=2/3(|x|)-4

OpenStudy (anonymous):

http://tinyurl.com/o4gtvty g(x) = ƒ(3x) http://tinyurl.com/lz8ovo8 g(x) = ƒ(2x) Those are my choices